Find the slope -6x - 9y = 12
Found 2 solutions by fractalier, checkley71:
Answer by fractalier(2101)

If we can get the line into y = mx + b form, we have the slope m.
We need to solve for y in
-6x - 9y = 12
-9y = 6x + 12
y = (-2/3)x - 4/3
So our slope is -2/3.

Answer by checkley71(8403)

-9y=6x+12 or y=-6x/9-12/9 or y=-2x/3-4/3 or the slope is -2/3